# On the cardinality of measures of maximal relative entropy for smooth skew products

### v2

@@ 3. Topological Markov shifts
$\Sigma$ $=\overline{\{\sigma^{i}(\underline{y})\}_{i\in\mathbb{Z}}}=\overline{\{\sigma^{i}\circ\mathrm{proj}_{\Sigma}(\underline{x})\}_{i\in\mathbb{Z}}}=\overline{\{\mathrm{proj}_{\Sigma}\circ\sigma^{i}_{s}(\underline{x})\}_{i\in\mathbb{Z}}}=\mathrm{proj}_{\Sigma}\left(\overline{\{\sigma^{i}_{s}(\underline{x})\}_{i\in\mathbb{Z}}}\right)$
$=\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})\subset\mathrm{proj}_{\Sigma}(\Sigma_{s})=\Sigma.$
In this way $\Sigma=\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})$. As in (7), we have that $\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s}^{\prime})$ is an subshift of finite type. Moreover, since $\Sigma^{\prime}_{s}$ is irreducible, it is clear that $\Sigma_{b}^{\prime}=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s}^{\prime})$ will also be irreducible, which concludes the proof of the theorem.
4. Measures of Relative Maximal Entropy on Topological Markov Shifts
@@
    $\pi:X\to Y$ a surjective $1$-block factor map;
@@
In the case that $X$ is a countable topological Markov shift and $\mu$ is a $\sigma_{X}$-invariant probability measure, we denote $h_{\mathrm{top}}(\sigma_{X})$ as its Gurevich topological entropy and $h_{\mu}(\sigma_{X})$ its metric entropy, for details see [46] (see also, [24, 23]).
Definition 4.2 ([44]).
@@
Given a topological Markov shift $(X,\sigma)$, we denote as $\mathcal{A}_{X}$ its alphabet. Also, given $\underline{x}\in X$ we denote as $x_{i}$ its $i$-th coordinate. Let $a\in A_{X}$, we define the $1$-cylinder at coordinate $k$ as $a]_{k}:=\{\underline{x}\in X;x_{k}=a\}.$
@@
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich topological entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, if $\mu_{1}$ and $\mu_{2}$ are ergodic measures of maximal $\nu$-relative entropy then $\mu_{1}$ and $\mu_{2}$ are $\nu$-relatively orthogonal.
@@
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich topological entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, $(X,\sigma_{X})$ can only admit countably many ergodic measures of maximal $\nu$-relative entropy.
@@
From now on, we fix $X$ as an irreducible countable-state topological Markov shift, $Y$ as an irreducible subshift of finite type, and $\pi:X\to Y$ a continuous surjective $1$-block factor map.
In the following, we will construct a subshift of finite type $X_{n}$ which, in some sense, approximates $X$. Since $\pi$ is a $1$-block factor map we have that $\gamma:=\{\pi^{-1}([i]_{0})\in\mathcal{B}(X);i\in\mathcal{A}_{Y}\}$ is a partition of $X$ such that each $\gamma$ is a union of $1$-cylinders at coordinate 0. For each $n\in\mathbb{N}$ we define the partition of $X$ as
$\alpha_{n}:=\left\{[a_{1}],[a_{2}],\ldots,[a_{n}],X\setminus\left(\bigcup_{i=1}^{n}[a_{i}]\right)\right\}\bigvee\gamma,$
@@
Using the partition $\alpha_{n}$, we define a subshift of finite type $X_{n}$ as follows. Let $\mathscr{G}_{n}=(\mathscr{V}_{n},\mathscr{E}_{n})$ be the directed graph with vertex set $\mathscr{V}_{n}:=\alpha_{n}$ and edge set
$\mathscr{E}_{n}:=\bigl\{A\to A^{\prime};\,A,A^{\prime}\in\alpha_{n}\ \text{and }A\cap\sigma_{X}^{-1}(A^{\prime})\neq\emptyset\bigr\}.$
@@
$(\mathrm{proj}_{X_{n}}(\underline{x}))_{k}=A\ \text{whenever}\ \sigma_{X}^{k}(x)\in A,\ A\in\alpha_{n},\ k\in\mathbb{Z}.$
Also, since the partition $\alpha_{n}$ refines the partition $\gamma$, for each $A\in\alpha_{n}$ there exists a unique $i(A)\in\{[i_{0}];i_{0}\in\mathcal{A}_{Y}\}$ such that $\pi(A)\subset[i(A)]_{0}\subset Y.$ This labels each atom $A\in\alpha_{n}$ by $i(A)$ and induces a $1$-block factor map $\pi_{n}:X_{n}\to Y$ (induced by $\pi$ defined by
@@
Given an invariant measure $\mu$ of $(X,\sigma)$ such that $\pi_{*}\mu=\nu$, observe that $\mu_{n}:=(\mathrm{proj}_{X_{n}})_{*}\mu$ satisfies $(\pi_{n})_{*}\mu_{n}=\nu$. From the Kolmogorov-Sinai theorem (see [51, Theorem 4.9]) we have that
@@
where $\beta_{n}$ is the partition of $X_{n}$ 11-cylinders at coordinate 0. Since $\mathcal{B}(X)=\mathcal{B}\left(\bigcup_{n}\sigma(\alpha_{n})\right)$ from [51, Theorem 4.14]
@@ 4.1. Petersen-Quas-Shin construction
Fix $n\in\mathbb{N}$. Let $\mu_{n}^{1},\mu_{n}^{2}$ be two distinct ergodic $\sigma_{X_{n}}$-invariant measures on $X_{n}$ with $(\pi_{n})_{*}\mu_{n}^{1}=(\pi_{n})_{*}\mu_{n}^{2}=\nu,$ and assume $\mu_{n}^{1}$ and $\mu_{n}^{2}$ are not $\nu$-relatively orthogonal (Definition 4.2). Following [44], there exists an invariant measure $\mu_{n}^{3}$ on $X_{n}$ with $(\pi_{n})_{*}\mu_{n}^{3}=\nu$ and
@@
We sketch the construction of $\mu_{n}^{3}$. Diagram 2 summarises all the factor maps used in this construction in this section. Let $Z_{n}:=X_{n}\times X_{n}\times\{0,1\}^{\mathbb{Z}}$ and define the map $T:=\sigma_{X_{n}}\times\sigma_{X_{n}}\times\sigma_{\mathrm{Ber}}$, where $\sigma_{\mathrm{Ber}}$ is the shift on $\{0,1\}^{\mathbb{Z}}$. Let $\zeta$ be the (1/2,1/2)(1/2,1/2)–Bernoulli measure on $\{0,1\}^{\mathbb{Z}}$. Also, set $\widehat{\mu}_{n}:=\mu_{n}^{1}\otimes_{\nu}\mu_{n}^{2}$ and $\eta_{n}=\widehat{\mu}_{n}\otimes\zeta$. Observe that $\eta_{n}$ is TT–invariant.
@@
    $N_{k}(u,v):=\sup\{m<k:u_{m}=v_{m}\},\,\text{with the convention }\sup\emptyset=-\infty.$
    Let $r_{-\infty}:\{0,1\}^{\mathbb{Z}}\to\{0,1\}$ be a random variable taking the values 0 or $1$ according to the law $\zeta$. Define
    $\bigl(\mathrm{proj}_{X_{n}}^{3}(u,v,r)\bigr)_{k}:=\begin{cases}u_{k},&\text{if }r_{N_{k}(u,v)}=0,\\[2.0pt] v_{k},&\text{if }r_{N_{k}(u,v)}=1.\end{cases}$
@@
Now, we quote a result from [44] which is contained in the proof of Theorem 1 (see [44, Equation (45)]).
Theorem 4.6 ([44]).
Under the above assumptions and notations, for each $n\in\mathbb{N}$
$h_{\mu_{n}^{3}}(\sigma_{X_{n}})-h_{\nu}(\sigma_{Y})\geq\frac{h_{\mu_{n}^{1}}(\sigma_{X_{n}})+h_{\mu_{n}^{2}}(\sigma_{X_{n}})}{2}-h_{\nu}(\sigma_{Y})+\int_{S_{n}}\Xi_{n}(z)\,\eta_{n}(\mathrm{d}z),$
@@
$S_{n}:=\left\{z\in Z_{n};\,(\mathrm{proj}_{X_{n}}^{1}(z))_{-1}=(\mathrm{proj}_{X_{n}}^{2}(z))_{-1}\,\right\}.$
The integrand $\Xi_{n}(z)$ is given by
$\Xi_{n}(z)=\sum_{j\in\mathcal{A}_{X_{n}}}\psi\left(\frac{A_{n}^{j}(z)+B_{n}^{j}(z)}{2}\right)-\frac{1}{2}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\big(A_{n}^{j}(z)\big)-\frac{1}{2}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\big(B_{n}^{j}(z)\big),$
@@
$A_{n}^{j}(z):=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\,\big|\,\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\big)(z),\ B_{n}^{j}(z):=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\,\big|\,\mathcal{B}_{2,n}^{-}\vee\mathcal{B}_{0,n}\big)(z),$
@@
$\mathcal{B}_{1,n}^{-}:=\sigma\left(\,\{z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{1}(z))_{-m}=a\,\};\,a\in\mathcal{A}_{X_{n}},\ m\in\mathbb{N}\right),$
$\mathcal{B}_{2,n}^{-}:=\sigma\left(\,\{z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{2}(z))_{-m}=a\,\};\,a\in\mathcal{A}_{X_{n}},\ m\in\mathbb{N}\right),$

### v3

@@ 3. Topological Markov shifts
$G^{0}:=\left\{\underline{Z}\in\Sigma:\;\underline{R^{0}}\in\bigcap_{n\geq 1}\overline{\bigcup_{k\geq n}\sigma^{k}(\underline{Z})}\right\}\ \text{and }G^{1}:=\left\{\underline{Z}\in\Sigma:\;\underline{R^{1}}\in\bigcap_{n\geq 1}\overline{\bigcup_{k\geq n}\sigma^{-k}(\underline{Z})}\right\}$
are dense $G_{\delta}$ subsets of $\Sigma$. By the Baire category theorem, $G:=G^{0}\cap G^{1}$ is dense in $\Sigma$. Let $\underline{R}$ be an element of $G$.
Observe that the set $\bigcap_{n\geq 1}\overline{\bigcup_{k\geq n}\sigma^{k}(\underline{R})}$ is closed, forward $\sigma$-invariant, and it contains $\underline{R^{0}}$. Hence
$\bigcap_{n\geq 1}\overline{\bigcup_{k\geq n}\sigma^{k}(\underline{R})}=\Sigma.$
Applying the same argument to $\sigma^{-1}$ we obtain that
$\bigcap_{n\geq 1}\overline{\bigcup_{k\geq n}\sigma^{-k}(\underline{R})}=\Sigma.$ (9)
Therefore every point of $\Sigma$ is an accumulation point of both the forward and the backward $\sigma$-orbit of $\underline{R}$ via $\sigma$.
Let $(\omega,x):=\pi(\underline{R})\in\mathcal{X}\subset\Omega\times M$. We claim that $\omega\in\Omega\setminus\left(\bigcup_{i\in\mathbb{Z}}\theta^{i}(\partial\mathscr{P})\right).$ Assume, for a contradiction, that the claim is false. Then, there exists $i\in\mathbb{Z}$ with
$\theta^{i}(\omega)\in\partial\mathscr{P}=\partial^{s}\mathscr{P}\cup\partial^{u}\mathscr{P}.$
We treat the case $\theta^{i}(\omega)\in\partial^{s}\mathscr{P}$; the unstable boundary case is analogous with backward time. Since the Borel homoclinic class $\mathcal{X}$ admits a $\Theta$-invariant ergodic measure projecting to the full-supported ergodic measure $\mathbb{P}$, by Theorem 3.5 (2) there exists $\underline{R}^{*}\in\Sigma$ with
$\pi(\underline{R}^{*})=(\omega^{*},x^{*})\ \text{and}\ \omega^{*}\in\Omega\setminus\partial^{s}\mathscr{P}.$
Because the forward orbit of $\underline{R}$ is dense, there is an increasing sequence $k_{j}\to\infty$ such that $\sigma^{k_{j}}(\underline{R})\longrightarrow\underline{R^{*}}$ as $j\to\infty$. By Theorem 3.5 (1), we have
$\mathrm{dist}_{\Omega}\big(\theta^{k_{j}}(\omega),\omega^{*}\big)$ $\leq\mathrm{dist}_{\Omega\times M}\big(\Theta^{k_{j}}(\omega,x),(\omega^{*},x^{*})\big)=\mathrm{dist}_{\Omega\times M}\big(\Theta^{k_{j}}\pi(\underline{R}),\pi(\underline{R^{*}})\big)$
$\leq C\,\mathrm{dist}_{\Sigma}\big(\sigma^{k_{j}}(\underline{R}),\underline{R^{*}}\big)^{\eta}\to 0\ \text{as }j\to\infty$ (10)
for suitable constants $C,\eta>0$.
On the other hand, by [BowenBook, Prop. 3.5],
$\theta(\partial^{s}\mathscr{P})\subset\partial^{s}\mathscr{P}\ \text{and}\ \theta^{-1}(\partial^{u}\mathscr{P})\subset\partial^{u}\mathscr{P}.$
Hence from $\theta^{i}(\omega)\in\partial^{s}\mathscr{P}$ we get $\theta^{i+n}(\omega)\in\partial^{s}\mathscr{P}$ for all $n\geq 0$. In particular, for all $j$ large enough so that $k_{j}\geq i$, $\theta^{k_{j}}(\omega)\in\partial^{s}\mathscr{P}.$ Since $\partial^{s}\mathscr{P}$ is closed and $\omega^{*}\notin\partial^{s}\mathscr{P}$, the distance $\delta:=\mathrm{dist}_{\Omega}\big(\omega^{*},\partial^{s}\mathscr{P}\big)>0,$ and therefore $\mathrm{dist}_{\Omega}\big(\theta^{k_{j}}(\omega),\omega^{*}\big)\geq\delta$ for all such $j$ large enough. This contradicts (10). The case $\theta^{i}(\omega)\in\partial^{u}\mathscr{P}$ is analogous. Hence $\omega\notin\bigcup_{i\in\mathbb{Z}}\theta^{i}(\partial\mathscr{P})$, proving the claim.
Using that $\omega\notin\bigcup_{i\in\mathbb{Z}}\theta^{i}(\partial\mathscr{P})$, from the proof of Step 3, there exists $\underline{(P,R)}\in\Sigma_{s}$ such that $\pi_{s}(\underline{(P,R)})=(\omega,x)$ and $\mathrm{proj}_{\Sigma}(\underline{(P,R)})=\underline{R}$. Let $\mathscr{G}^{\prime}_{s}\subset\mathscr{G}_{s}$ be the maximal irreducible component of $\mathscr{G}_{s}$ containing $(P_{0},R_{0})$ (the 0-th element of $\underline{(P,R)}$) and define $\Sigma^{\prime}_{s}:=\Sigma(\mathscr{G}^{\prime}_{s})\subset\Sigma(\mathscr{G}_{s})=\Sigma_{s}.$ Observe that $\underline{(P,R)}\in\Sigma_{s}^{\prime}$ and therefore $\sigma^{i}_{s}(\underline{(P,R)})\in\Sigma_{s}^{\prime}$ for each $i\in\mathbb{Z}.$
We claim that $\operatorname{proj}_{\Sigma}(\Sigma^{\prime}_{s})=\Sigma$. First, density: let $U\subset\Sigma$ be nonempty open. Since the orbit of $\underline{R}$ is dense, there exists $i\in\mathbb{Z}$ with $\sigma^{i}(\underline{R})\in U$. By construction, $\sigma^{i}(\underline{R})=\operatorname{proj}_{\Sigma}\left(\sigma_{s}^{i}(\underline{(P,R)})\right)$ with $\sigma_{s}^{i}(\underline{(P,R)})\in\Sigma^{\prime}_{s}$, hence $U\cap\mathrm{proj}_{\Sigma}(\Sigma^{\prime}_{s})\neq\varnothing$. Thus $\mathrm{proj}_{\Sigma}(\Sigma^{\prime}_{s})$ is dense.
To see that $\operatorname{proj}_{\Sigma}(\Sigma^{\prime}_{s})$ is closed, let $\{\underline{R}^{(m)}\}_{m\geq 1}\subset\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})$ with $\underline{R}^{(m)}\to\underline{R}$ in $\Sigma$ as $m\to\infty$. For each $m$ choose a lift $\underline{Y}^{(m)}=\{(P^{(m)}_{j},R^{(m)}_{j})\}_{j\in\mathbb{Z}}\in\Sigma_{s}^{\prime}$ such that $\mathrm{proj}_{\Sigma}(\underline{Y}^{(m)})=\underline{R}^{(m)}$ (so $R^{(m)}_{j}=\underline{R}^{(m)}_{j}$ for all $j$). Since $\Sigma_{b}$ is a subshift of finite type, it is compact. Therefore, passing through a subsequence if necessary, we can assume that $\underline{P}^{(m)}:=\{P^{(m)}_{j}\}_{j\in\mathbb{Z}}\to\underline{P}=\{P_{j}\}_{j\in\mathbb{Z}}$ in $\Sigma_{b}$ as $m\to\infty$. Because $\underline{R}^{(m)}\to\underline{R}$, for each fixed $j$ we have $R^{(m)}_{j}=R_{j}$ for all large $m$, and therefore,
$\underline{Y}^{(m)}=\{(P^{(m)}_{j},R^{(m)}_{j})\}_{j\in\mathbb{Z}}\longrightarrow\underline{Y}:=\{(P_{j},R_{j})\}_{j\in\mathbb{Z}}.$
The space $\Sigma_{s}^{\prime}$ is a countable-state Markov shift (edge shift), hence closed; thus $\underline{Y}\in\Sigma_{s}^{\prime}$. Finally, $\mathrm{proj}_{\Sigma}(\underline{Y})=\underline{R}$, so $\underline{R}\in\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})$. This proves that $\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})$ is closed.
In this way $\Sigma=\mathrm{proj}_{\Sigma}(\Sigma_{s}^{\prime})$. As in Step 5, the same proof yields that $\widehat{\mathbb{P}}[\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})]=1.$
4. Measures of maximal relative entropy on topological Markov shifts
@@
    $\pi:X\to Y$ a $1$-block map, $\nu(\pi(X))=1$ and $\mathrm{supp}(\nu)=Y$;
@@
In the case that $(X,\sigma_{X})$ is a topological Markov shift and $\mu$ is a $\sigma_{X}$-invariant probability measure, we denote $h_{\mathrm{top}}(\sigma_{X})$ as its Gurevich topological entropy and $h_{\mu}(\sigma_{X})$ its metric entropy, for details see [SarigThermo] (see also, [Gurevich1, Gurevich2]).
Definition 4.2 ([Quas]).
@@
Given a topological Markov shift $(X,\sigma)$, we denote as $\mathcal{A}_{X}$ its alphabet. Also, given $x\in X$ we denote as $x_{i}$ its $i$-th coordinate. Let $a\in A_{X}$, we define the $1$-cylinder at coordinate $k$ as $a]_{k}:=\{x\in X;x_{k}=a\}.$
@@
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich topological entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a $1$-block map such that $\nu(\pi(X))=1$ and $\mathrm{supp}(\nu)=Y$. Then, if $\mu_{1}$ and $\mu_{2}$ are ergodic measures of maximal $\nu$-relative entropy then $\mu_{1}$ and $\mu_{2}$ are $\nu$-relatively orthogonal.
@@
Let $(X,\sigma_{X})$ be an irreducible topological Markov shift with finite Gurevich topological entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a $1$-block map such that $\nu(\pi(X))=1$ and $\mathrm{supp}(\nu)=Y$. Then, $(X,\sigma_{X})$ can only admit countably many ergodic measures of maximal $\nu$-relative entropy.
@@
From now on, we fix $(X,\sigma_{X})$ as an irreducible topological Markov shift with countable alphabet, $(Y,\sigma_{Y})$ as an irreducible subshift of finite type, and $\pi:X\to Y$ a continuous $1$-block map such that $\nu(\pi(X))=1$.
In the following, we will construct a subshift of finite type $(X_{n},\sigma_{X_{n}})$ which, in some sense, approximates $(X,\sigma_{X})$. Since $\pi$ is a $1$-block map we have that $\gamma:=\{\pi^{-1}([i]_{0})\in\mathcal{B}(X);i\in\mathcal{A}_{Y}\}$ is a partition of $X$ such that each $\gamma$ is a union of $1$-cylinders at coordinate 0. For each $n\in\mathbb{N}$ we define the partition of $X$ as
$\alpha_{n}:=\left\{[a_{1}]_{0},[a_{2}]_{0},\ldots,[a_{n}]_{0},X\setminus\left(\bigcup_{i=1}^{n}[a_{i}]\right)\right\}\bigvee\gamma,$
@@
Using the partition $\alpha_{n}$, we define a subshift of finite type $(X_{n},\sigma_{X_{n}})$ as follows. Let $\mathscr{G}_{n}=(\mathscr{V}_{n},\mathscr{E}_{n})$ be the directed graph with vertex set $\mathscr{V}_{n}:=\alpha_{n}$ and edge set
$\mathscr{E}_{n}:=\bigl\{A\to A^{\prime};\,A,A^{\prime}\in\alpha_{n}\ \text{and }A\cap\sigma_{X}^{-1}(A^{\prime})\neq\varnothing\bigr\}.$
@@
$(\mathrm{proj}_{X_{n}}(x))_{k}=A\ \text{whenever}\ \sigma_{X}^{k}(x)\in A,\ A\in\alpha_{n},\ k\in\mathbb{Z}.$
Also, since the partition $\alpha_{n}$ refines the partition $\gamma$, for each $A\in\alpha_{n}$ there exists a unique $i(A)\in\{[i_{0}];i_{0}\in\mathcal{A}_{Y}\}$ such that $\pi(A)\subset[i(A)]_{0}\subset Y.$ This labels each atom $A\in\alpha_{n}$ by $i(A)$ and induces a $1$-block map $\pi_{n}:X_{n}\to Y$ (induced by $\pi$ defined by
@@
Let $\mu$ be a $\sigma$-invariant measure on $X$ with $\pi_{*}\mu=\nu$, and for each $n\in\mathbb{N}$ set $\mu_{n}=(\mathrm{proj}_{X_{n}})_{*}\mu$. Since $\pi=\pi_{n}\circ\mathrm{proj}_{X_{n}}$, we have
$(\pi_{n})_{*}\mu_{n}=(\pi_{n}\circ\mathrm{proj}_{X_{n}})_{*}\mu=\pi_{*}\mu=\nu.$
Hence $\nu$ is supported on $\pi_{n}(X_{n})$, i.e. $\nu\big(Y\setminus\pi_{n}(X_{n})\big)=0$. Because $\pi_{n}(X_{n})$ is closed in $Y$ and $\mathrm{supp}\,(\nu)=Y$, it follows that $\pi_{n}(X_{n})=Y$. Therefore $\pi_{n}:X_{n}\to Y$ is surjective for every $n\in\mathbb{N}$.
From the Kolmogorov-Sinai theorem (see [WaltersBook, Theorem 4.9]) we have that
@@
where $\beta_{n}$ is the partition of $X_{n}$ 11-cylinders at coordinate 0. Since, $\alpha_{1}\preccurlyeq\alpha_{2}\preccurlyeq\ldots$ and $\mathcal{B}(X)=\mathcal{B}\left(\bigvee_{n\in\mathbb{N}}\alpha_{n}\right)$ from [WaltersBook, Theorem 4.14]
@@
With a slight abuse of notation, we write $a_{1},\ldots,a_{n}\in\mathcal{A}_{X_{n}}$ whenever the context is clear.
@@ 4.1. Petersen-Quas-Shin construction
Select and fix $a\in A_{X}$. Choose $n$ large enough that $a]_{0}\in\alpha_{n}$ (so $a\in A_{X_{n}}$). Let $\mu_{n}^{1},\mu_{n}^{2}$ be two distinct ergodic $\sigma_{X_{n}}$-invariant measures on $X_{n}$ with $(\pi_{n})_{*}\mu_{n}^{1}=(\pi_{n})_{*}\mu_{n}^{2}=\nu,$ and assume $\mu_{n}^{1}$ and $\mu_{n}^{2}$ are not $\nu$-relatively orthogonal (Definition 4.2). Following [Quas], there exists an invariant measure $\mu_{n}^{3}$ of $(X_{n},\sigma_{X_{n}})$ with $(\pi_{n})_{*}\mu_{n}^{3}=\nu$ and
@@
We sketch the construction of $\mu_{n}^{3}$. Diagram 2 summarises all semi-conjugacies used in this construction in this section. Recall, from the previous section, that $\pi_{n}:X\to Y$ is surjective for every $n\in\mathbb{N}$. Let $Z_{n}:=X_{n}\times X_{n}\times\{0,1\}^{\mathbb{Z}}$ and define the map $T:=\sigma_{X_{n}}\times\sigma_{X_{n}}\times\sigma_{\mathrm{Ber}}$, where $\sigma_{\mathrm{Ber}}$ is the shift on $\{0,1\}^{\mathbb{Z}}$. Let $\zeta$ be the (1/2,1/2)(1/2,1/2)–Bernoulli measure on $\{0,1\}^{\mathbb{Z}}$. Also, set $\widehat{\mu}_{n}:=\mu_{n}^{1}\otimes_{\nu}\mu_{n}^{2}$ and $\eta_{n}=\widehat{\mu}_{n}\otimes\zeta$. Observe that $\eta_{n}$ is TT–invariant.
@@
    $N_{k}^{n}(u,v):=\sup\{m<k;u_{m}=v_{m}=a\},\,\text{with the convention }\sup\varnothing=-\infty.$ (11)
    $r\in\{0,1\}^{\mathbb{Z}}$, $r_{n}$ denotes its $n$-th coordinate, and let
    $r\in\{0,1\}^{\mathbb{Z}}\mapsto r_{-\infty}\in\{0,1\}$ (12)
[…]
    $\bigl(\mathrm{proj}^{3}_{X_{n}}(u,v,r)\bigr)_{k}:=\begin{cases}u_{k},&\text{if }r_{N^{n}_{k}(u,v)}=0,\\ v_{k},&\text{if }r_{N^{n}_{k}(u,v)}=1.\end{cases}$
@@
Now, we quote a result from [Quas] which is contained in the proof of Theorem 1 (see [Quas, Equation (45)]).
Theorem 4.6 ([Quas]).
Under the above assumptions and notations, for any $j_{*}\in\mathcal{A}_{X_{n}}$ for each $n\in\mathbb{N}$
$h_{\mu_{n}^{3}}(\sigma_{X_{n}}\mid\nu)\geq\frac{h_{\mu_{n}^{1}}(\sigma_{X_{n}}\mid\nu)+h_{\mu_{n}^{2}}(\sigma_{X_{n}}\mid\nu)}{2}+\int_{S_{n}^{(a)}}\Xi_{n}^{j_{*}}(z)\,\eta_{n}(\mathrm{d}z),$
@@
$S_{n}^{(a)}:=\left\{z\in Z_{n};\,(\mathrm{proj}_{X_{n}}^{1}(z))_{-1}=(\mathrm{proj}_{X_{n}}^{2}(z))_{-1}=a\,\right\}.$
The integrand $\Xi_{n}^{j_{*}}(z)$ is given by
$\Xi_{n}^{j_{*}}(z)=\psi\left(\frac{A_{n}^{j_{*}}(z)+B_{n}^{j_{*}}(z)}{2}\right)-\frac{1}{2}\psi\big(A_{n}^{j_{*}}(z)\big)-\frac{1}{2}\psi\big(B_{n}^{j_{*}}(z)\big),$
@@
$A_{n}^{j_{*}}(z):=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j_{*}]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\,\big|\,\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\big)(z),\ B_{n}^{j_{*}}(z):=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j_{*}]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\,\big|\,\mathcal{B}_{2,n}^{-}\vee\mathcal{B}_{0,n}\big)(z),$
@@
$\mathcal{B}_{1,n}^{-}:=\sigma\left(\,\{z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{1}(z))_{-m}=j\,\};\,j\in\mathcal{A}_{X_{n}},\ m\in\mathbb{N}\right),$
$\mathcal{B}_{2,n}^{-}:=\sigma\left(\,\{z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{2}(z))_{-m}=j\,\};\,j\in\mathcal{A}_{X_{n}},\ m\in\mathbb{N}\right),$